### v1

# Securing Automated Insulin Delivery Systems: A Review of Security Threats and Protectives Strategies

Abstract
Automated insulin delivery (AID) systems have emerged as a significant technological advancement in diabetes care. These systems integrate a continuous glucose monitor, an insulin pump, and control algorithms to automate insulin delivery, reducing the burden of self-management and offering enhanced glucose control. However, the increasing reliance on wireless connectivity and software control has exposed AID systems to critical security risks that could result in life-threatening treatment errors. This review first presents a comprehensive examination of the security landscape, covering technical vulnerabilities, legal frameworks, and commercial product considerations, and an analysis of existing research on attack vectors, defence mechanisms, as well as evaluation methods and resources for AID systems. Despite recent advancements, several open challenges remain in achieving secure AID systems, particularly in standardising security evaluation frameworks and developing comprehensive, lightweight, and adaptive defence strategies. As one of the most widely adopted and extensively studied physiologic closed-loop control systems, this review serves as a valuable reference for understanding security challenges and solutions applicable to analogous medical systems.
keywords:
automated insulin delivery systems, artificial pancreas systems , security , cybersecurity , attack vectors, defence , physiologic closed-loop control system
[…]
[label1]organization=College of Computing and Data Science, Nanyang Technological University,addressline=50 Nanyang Ave, postcode=639798, country=Singapore
1 Introduction
The management of Type 1 diabetes (T1D), a chronic condition where pancreatic beta cells produce insufficient insulin for regulating blood glucose (BG) levels, remains a significant challenge for millions of individuals worldwide [1]. As there is no realistic biological cure in the near future, the treatment primarily involves intensive manual insulin therapy to manage BG levels in the range to prevent hyperglycemia or hypoglycemia, both of which can lead to severe, life-threatening complications [2]. However, this treatment regimen requires frequent BG monitoring and insulin adjustment, which are time-consuming and error-prone. To address these challenges, automated insulin delivery (AID) systems, also known as artificial pancreas (AP) systems, have emerged as a significant technological advancement in diabetes care [3].
AID system is a wearable physiological closed-loop control system that integrates a continuous glucose monitor (CGM), an insulin pump, and control algorithms to modulate insulin delivery based on continuous glucose measurements. This system not only alleviates the burden of self-management but also significantly improves clinical outcomes [4]. The demand for AID systems surged by advanced technology and the growing global prevalence of diabetes [5][6]. In 2023, the global AID systems market reached US\$ 2,338.5 million and is projected to expand to US\$ 5,800.2 million by 2032 [7]. However, the complex interconnectivity, increasing reliance on automated control, and wireless connectivity for wearable convenience in AID systems introduce significant risks that can lead to real-time treatment errors with potentially life-threatening consequences.
Threats to the reliability and safety of AID systems primarily arise from two sources: unintended failures (i.e., device limitations and malfunctions) and malicious attacks. Unintended failures stem from device physiological, technological, and behavioural limitations [3], as well as system malfunctions [8]. For instance, physiological constraints on delayed insulin absorption and postprandial rapid glucose fluctuations can render ineffective insulin delivery of AID systems. Technological limitations, including inaccurate glucose readings due to noise spikes, signal dropouts, miscalibration, or pressure applied to the sensor site (e.g., lying on the sensor while sleeping), can result in inappropriate insulin dosing. Additionally, behavioural limitations, such as improper patient operation, and malfunctions in critical components like infusion site failures and cartridge errors, can also lead to AID systems failures.
Malicious attacks, on the other hand, involve adversaries exploiting vulnerabilities in AID systems to intercept patient data or impact insulin delivery. These cybersecurity threats have escalated with the adoption of wireless communication in AID systems [9]. Examples include unauthorised remote control [10] and insulin dose manipulation attacks [11]. Early experiments have shown the feasibility of cyberattacks through unencrypted wireless channels to deliver lethal insulin doses from at least 20 meters away [12][13]. The U.S. Food & Drug Association (FDA) issued warnings in 2019 regarding AID systems cybersecurity weaknesses [14], and commercial AID systems products have been recalled due to security concerns [15].
Review Paper Focus Discussion on Unintended Failures Discussion on Malicious Attacks Analysis of Attack Vectors Evaluation of Defence Mechanism Klonoff, 2015 [16] Cybersecurity for Connected Diabetes Devices $\times$ $\checkmark$ $\times$ (no literature review) O’Keeffe et al,. 2015 [17] Cybersecurity in AID systems $\times$ $\checkmark$ $\times$ (no literature review) Blauw et al., 2016 [18] Safety and design requirements of the AID systems $\checkmark$ $\times$ $\times$ (For unintended failures) Ramkissoon et al., 2017 [19] Specific hazards applicable to the AID systems $\checkmark$ $\times$ $\times$ (For unintended failures) Thompson et al., 2018 [20] Safety issues in outpatient settings $\checkmark$ $\checkmark$ $\times$ $\times$ Quintal er al., 2019 [21] Ethical issues associated with AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Singh and Rahman, 2020 [22] The security of Open AID systems system $\times$ $\checkmark$ (no literature review) (no literature review) Sherr et al., 2022 [3] Current landscape and safety challenges of AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Wirth, 2022 [23] The states and tendency of cybersecurity in the diabetes ecosystem $\times$ $\checkmark$ $\times$ $\times$ Nazzal et al., 2024 [24] Software safety and security verification and validation on AID systems $\times$ $\times$ $\times$ (Formal Verification methods) Kapadia, 2024 [25] Limitations and malfunctions $\checkmark$ $\times$ $\times$ $\times$ This paper The cybersecurity status, attack vectors, and proposed defence strategies of AID systems $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of review papers on the security of AID systems in Last 10 Years.
Strengthening AID systems security requires a thorough understanding of existing and emerging threats, a critical evaluation of defence mechanisms, and the identification of gaps that must be addressed. As summarized in Table 1, most review papers published in the past decade have primarily focused on unintended failures [18] [19] [25] or discussed threats from malicious attacks without analysing specific attack vectors and defence mechanisms [26] [21] [3]. While several studies [16] [17] [27] [28] provided general overviews of cyber vulnerabilities and potential countermeasures, they lack in-depth examinations of related studies. Nazzal et al. [25] conducted a thorough review of software verification methods for AID systems; however, broader security defence strategies remain unexplored.
Besides, compared to most other safety-critical devices within the Internet of Medical Things (IoMT), which also face threats from malicious attacks, AID systems present unique security challenges: (1) the human-in-the-loop component, which introduces individual physiological variability and unpredictable external factors, especially in free-living settings, and (2) constrained power, computational, and memory resources, which limit the feasibility of typical resource-intensive security strategies. These challenges not only amplify the impact of security threats but also complicate the development of effective defence mechanisms. While existing IoMT security reviews [29][30][31] discussed AID systems cybersecurity, they fail to address its unique risks. Notably, no prior reviews comprehensively examine security threats and defence strategies for wearable physiologic closed-loop control systems.
To address this gap, this study systematically examines the security of AID systems, providing a detailed analysis of attack vectors, existing defence strategies, and potential areas for future research. We introduce a taxonomy that categorizes security threats and defence mechanisms in AID systems, as shown in Figure 2. Given the increasing adoption of AID systems and the growing body of research in this field, this review offers valuable insights for enhancing the security of similar medical systems.
The paper is organized as follows (see Figure 1). Section 3 explores AID systems cybersecurity from technical, legal, and commercial perspectives, establishing a foundational understanding of the field. Section 4 examines existing and emerging attack vectors, followed by Section 5, which reviews existing defence strategies. Section 6 discusses current evaluation methods and available research resources. Finally, Section 7 summarizes key findings and identifies research gaps, and Section 8 outlines future directions for strengthening AID systems security.
Figure 1: Overview of the paper.
Figure 2: Taxonomy of security threats and defence strategies in AID systems.
2 Methodology and Materials
To comprehensively assess the security landscape of AID systems, this review adopts a hybrid methodology, integrating elements of both survey and systematic review. The review begins with a broad investigation into AID systems cybersecurity, covering technical vulnerabilities, regulatory frameworks, and security measures in commercial products. The materials considered extend beyond peer-reviewed journals and conference papers to include diverse sources such as clinical studies, industry reports, regulatory standards, product manuals, and relevant online resources. Following this, the PRISMA framework [32] was employed for systematic analysis of attack vectors and defence mechanisms for AID systems security. To guide this analysis, the following research questions (RQs) were formulated:
  • 1.
    RQ1: What potential attack vectors and risks do AID systems face?
  • 2.
    RQ2: How do existing defence mechanisms address these risks?
  • 3.
    RQ3: How practical and robust are these solutions in protecting AID systems from cybersecurity threats?
A systematic literature search was conducted across multiple databases, including Scopus, Google Scholar, Web of Science, IEEE Xplore, and PubMed. Boolean operators and keyword variations were applied, as summarized in Table 2. The review focused on publications from 2010 to the collection date (February 11, 2025), with studies outside this range filtered out during the initial search.
Search Engine Keywords Google Scholar (intitle:”artificial pancreas” OR intitle:”closed-loop insulin pump” OR intitle:”automated insulin delivery” OR intitle:”hybrid closed loop systems”) AND (attack OR cyberattack OR hijack OR cybersecurity OR security OR ”intrusion detection” OR secure) Scopus TITLE-ABS-KEY ( (”artificial pancreas” OR ”closed-loop insulin pump” OR ”automated insulin delivery” OR ”hybrid closed loop systems”) AND (attack OR cyberattack OR hijack OR cybersecurity OR security OR ”intrusion detection” OR secure)) IEEE Xplore (”artificial pancreas” OR ”closed-loop insulin pump” OR ”automated insulin delivery” OR ”hybrid closed-loop systems”) AND (attack OR cyberattack OR hijack OR cybersecurity OR security OR ”intrusion detection” OR secure) Web of Science (”artificial pancreas” OR ”closed-loop insulin pump” OR ”automated insulin delivery” OR ”hybrid closed-loop systems”) AND (attack OR cyberattack OR hijack OR cybersecurity OR security OR ”intrusion detection” OR secure) PubMed (”artificial pancreas” OR ”closed-loop insulin pump” OR ”automated insulin delivery” OR ”hybrid closed-loop systems”) AND (attack OR cyberattack OR hijack OR cybersecurity OR security OR ”intrusion detection” or secure)
Table 2: Keywords used in search engines.
All retrieved papers underwent a screening process (PRISMA diagram in Figure 3) based on predefined inclusion and exclusion criteria.
Inclusion criteria:
  • 1.
    Peer-reviewed journal articles, conference papers, or book chapters;
  • 2.
    Published between 2010 and 2025;
  • 3.
    Studies proposing attack vectors, defence mechanisms, or risk evaluations for artificial pancreases
Exclusion criteria:
  • 1.
    Papers not written in English
  • 2.
    Review papers;
  • 3.
    Irrelevant articles;
  • 4.
    Duplicate articles
Figure 3: PRISMA diagram for the screening process.
From the initial database search, 53 relevant papers were identified. To ensure comprehensive coverage, backward and forward citation tracking was conducted, yielding 23 additional studies. These were initially missed due to: (1) Broad scope—studies focusing on medical devices, IoT-based healthcare systems, or cyber-physical systems that included AID systems as a case study; and (2) Terminology differences—studies referring to AID systems under alternative terms such as wireless insulin pump systems or sensor-augmented pumps.
Finally, 76 papers met the inclusion criteria for analysing attack vectors and defence mechanisms in AID systems. The publication trends over the past 15 years are visualized in Figure 4. Notably, a significant rise in proposed defence mechanisms has been observed in the past five years.
Figure 4: Publication trend of attack and defence literature.
3 Background
3.1 Technical Description and Vulnerabilities
Figure 5: An overview of a hybrid closed-loop insulin delivery system, illustrating core components, workflow and potential functional vulnerabilities. Dotted lines indicate wireless communication links between modules, while solid lines represent direct attachments to the patient’s body or integrated elements - adapted from [33].
AID combines a CGM, an insulin pump and a controller algorithm to calculate and deliver suitable insulin doses, helping to maintain blood glucose levels within the target range. Figure 5 shows the workflow of a hybrid closed-loop insulin delivery system, the paradigm common among current FDA-approved AID systems products. In this system, the CGM periodically (e.g., every 5 minutes) transmits the measured glucose levels to the control algorithm device, which calculates basal rates and corrective bolus doses of insulin, and then directs the subcutaneous insulin pump to deliver the specified dose into the patient body [34]. Unlike fully closed-loop systems that automate all insulin delivery, hybrid systems still require user input for meal and activity-related boluses [35]. Fully closed-loop systems, though promising, still face challenges in managing postprandial hyperglycemia and exercise-related glycemic fluctuations [36]. Despite different automation levels, both hybrid and fully closed-loop systems share core modules and rely on wireless connectivity and software control, exposing them to similar cybersecurity vulnerabilities.
In the context of AID systems, cybersecurity focuses on protecting digital assets, which encompass physiological data, treatment commands transmitted wirelessly between system components, the software controlling these processes, and the patient’s personal information stored within the device. Vulnerabilities exist both in wireless communication channels and individual system modules, as detailed in the following subsections.
3.1.1 Wireless communication
The communication between the three main components of AID systems primarily relies on wireless transmissions for wearable convenience. Safe information exchange requires secure, lightweight, and resilient communicational protocols, including proper authentication and encryption. However, as examined in [37], applied protocols often fall short of meeting safety requirements. For example, the widely used Bluetooth LE protocol is vulnerable to replay and man-in-the-middle attacks [38], allowing adversaries to eavesdrop on communications or inject faulty glucose data and erroneous control commands.
3.1.2 CGM
A CGM system continuously tracks blood glucose (BG) levels in the patient’s body and transmits the collected physiological data to the controller every few minutes. It consists of a disposable glucose sensor, placed under the skin to measure glucose levels, and a transmitter attached to the sensor that sends the data to the control algorithm platform. Accurate BG readings are crucial for safe insulin delivery, as any incorrect values can potentially result in harmful dosing decisions. However, a compromised CGM system may operate maliciously by providing inaccurate BG levels, either due to device tampering or data manipulation during wireless transmission [12][34][13].
3.1.3 Control Algorithms
The controller computes the recommended insulin doses and transmits infusion commands to the insulin pump. Its decision-making process is based on real-time sensor data, historical records, and estimated patient status (e.g., eventual BG, Insulin on Board (IOB)). The controller has a patient physiological model to estimate glucose-insulin dynamics in the patient’s body, and control algorithms to determine optimal insulin doses. These algorithms are typically deployed on smartphones, tablets, personal computers, or integrated with insulin pumps, which are often the platforms for user monitoring of device activities. Three main security threats are associated with the controller:
  • 1.
    Data Manipulation: The stored records, received sensor data, or meal and activities inputs could be tampered with or falsely registered, compromising the accuracy of the controller’s calculations [39][40]. Additionally, alerts on the display screen could be maliciously triggered or disabled, prompting users to turn off or ignore the security monitoring system, thus heightening system vulnerability.
  • 2.
    Compromising controller algorithms: The core logic and configuration of control algorithms may be altered, resulting in incorrect insulin dosing [41].
  • 3.
    Platform Vulnerabilities: The device running the control algorithm could be compromised by malware, such as viruses or worms, which may interfere with the proper functioning of the AID systems [42].
3.1.4 Subcutaneous Insulin Pump
A subcutaneous insulin pump includes a reservoir, which stores insulin and supplies it to the delivery mechanism, an infusion set consisting of a thin tube and a subcutaneous cannula for insulin delivery, and a microcontroller unit that orchestrates the pump’s functions by executing programmed commands based on input from the control platform. Security concerns associated with the pump arise from device malfunctions (e.g., infusion site failures)[43][44] and malicious control I/O operations of the pump[13].
3.2 Regulation Standards
Applied Area Regulation Standards Cybersecurity Note US Policy for Device Software Functions and Mobile Medical Applications Require adequate cybersecurity to secure the correct functioning of medical devices Provide specific cybersecurity considerations Section 3060 and 3060(a) of the 21st Century Cure Acts EO 14019 on America’s Supply Chains Require detailed technical cybersecurity measures used by the product and related services Not mandated for medical devices yet; could be in the future Guidance for FDA staff Cybersecurity in Medical Devices Provides expectations for inspection and approval, but no detailed federal guidance EU Medical Devices Regulation (MDR) No provisions specifically labeled ‘cybersecurity’ Highest priority MDR Annex1: General Safety and Performance Requirements (GSPR) Require that medical devices do not fail due to any type of adversary attacking them Imply minimum levels of encryption, access control, and physical cybersecurity Network and Information Systems (NIS 2) Directive Cybersecurity requirement of general critical IT infrastructure Integrated into Member States in various ways Cyber Resilience Act and the Cyber Solidarity Act Regulate the cybersecurity of any network-enabled devices Supposed to entail more resilience (greater recovery from failure) than MDR Security standards: ISO 14155:2011 Mostly developed by private actors and does not necessarily reflect best practices. Allow usage of the standard for a fee, which becomes a barrier to adoption General Data Protection Regulation (GDPR) Protecting the fundamental right to privacy Only considering cybersecurity requirements related to personal data
Table 3: Regulation standards of cybersecurity in medical devices.
The regulatory landscape governing AID systems cybersecurity is evolving, with growing recognition of safeguarding patient safety. Table 3 summarizes the core regulatory standards for AID systems cybersecurity in the United States (US) and European Union (EU). Given the global nature of medical device markets, the regulatory frameworks of many other countries closely align with these standards.
In the US, the FDA is responsible for overseeing medical devices’ compliance with regulations. Existing referable legal sources and guidances for FDA staff underscore the importance of cybersecurity as a key criterion in the FDA’s device inspection and approval processes [45]. In the Policy for Device Software Functions and Mobile Medical Applications, medical device manufacturers are required to implement adequate measures to protect against potential cyberattacks. Nonetheless, while the emphasis on cybersecurity is clear, there remains a notable absence of detailed, technical mandates similar to those applied in supply chain regulations. Given the increasing focus on medical device security, more specific technical requirements may be introduced in the future [46].
In the EU, the Medical Devices Regulation (MDR) serves as the main legislative framework, along with other cybersecurity standards. Although the MDR is positioned as the highest priority, it contains no provisions specifically labelled ‘cybersecurity’. Nevertheless, Annex 1 of the MDR, General Safety and Performance Requirements (GSPR) and other cybersecurity legislation provide a promising landscape regulation for cybersecurity in medical devices. For example, GSPR requires medical devices not to fail due to any type of security attack and that the necessary cybersecurity for functioning must be in place.
Generally, the EU and the US regulations mainly provide high-level principles rather than detailed technical specifications. While manufacturers are encouraged to adopt best practices like encryption, data access control, and software patching, there is no mandated standard for real-time monitoring or automatic threat response systems. A more specific review of cybersecurity in medical device regulation and its future consideration can be found in [46].
Beyond government regulations, a nonprofit organization named Diabetes Technology Society (DTS) proposed a Cybersecurity Standard for the Connected Diabetes Devices project [47]. This standard provides a standardized approach to thoroughly evaluating and confirming the conformity and confidence that the medical device can protect itself against applicable security threats, offering an example of how regulation could evolve toward more comprehensive and concrete evaluations.
3.3 Commercial Products
Within the regulation of existing cybersecurity frameworks, leading commercial AID system products demonstrate only basic cybersecurity capabilities, lacking comprehensive and robust protective measures, as outlined in Table 4. The DBLG1 System employed data encryption to secure wireless communication and Medtronic 780G includes a safety feature that pauses insulin delivery when the glucose level falls below a preset threshold, while others do not specify particular cybersecurity strategies.
Experiments [12][48] have demonstrated that at least two commercial products lack standard cryptographic mechanisms and the wireless communication protocol can be fully reverse-engineered, allowing eavesdropping, replay and command manipulation attacks to be launched. These vulnerabilities have led to commercial products being recalled [10] and warning of use [49]. Even devices with basic encryption remain susceptible to attacks that compromise integrity and availability. In the meantime, mechanisms hired to prevent patient harm are insufficient. Process-aware attacks [40] and personalized insulin dose manipulation attacks [50] can cause severe health consequences by subtly altering insulin delivery without injecting over-dose insulin to raise alarms.
Product Components Main Functions Cybersecurity CamAPS FX User’s smartphone (controller);
DANA/Ypsomed insulin pump;
Dexcom/Libre 3 CGM
Automated basal insulin delivery; Manual bolusing for meals User’s responsibility
Medtronic 780G MiniMed Insulin pump (controller integrated);
MiniMed CGM
Automated basal insulin delivery; Recommend precise bolus doses Mainly the user’s responsibility; Pause delivery when the glucose level is less than a preset level
Control IQ User’s smartphone (controller);
Tandem t: Slim insulin pump;
Dexcom G6/G7 CGM;
Automated basal insulin delivery; Automatic correction boluses delivery User’s responsibility
Omnipod 5 User’s smartphone (controller);
Omnipod insulin pumps;
Dexcom G6 CGM
Automatic basal insulin adjustments User’s responsibility
DBLG1 System DBLG1 controller;
Kaleido pump;
Dexcom G6 CGM
Automated basal insulin delivery; Automatic correction boluses delivery Mainly the user’s responsibility; Data encrypted
Table 4: Cybersecurity in representative AID system products.
Currently, according to product manuals, the responsibility for detecting and responding to security threats is placed on the patients themselves. However, this expectation is often impractical in real-world scenarios, particularly during sleep or daily activities. Additionally, the absence of a CGM display further limits the user’s ability to promptly detect deviations in BG readings from the control platform. Under these circumstances, AID system failures remain prevalent. A thesis [25] listed 24 documented device recalls and failures associated with a single AID systems manufacturer up until 2023 (Table 05). Despite evidence of their effectiveness in diabetes management, a cross-sectional survey [51] indicates that concerns over privacy and reliability may hinder AID systems adoption.
Given these challenges, it is imperative for both manufacturers and regulators to establish comprehensive and specific cybersecurity standards. Ensuring that AID systems devices can timely detect and mitigate security threats — without relying solely on user intervention—will be essential for enhancing system reliability, fostering user trust, and fully realizing the potential of AID systems in diabetes care.
4 Attack Vectors in AID systems
Building on the discussion of cybersecurity vulnerabilities in AID systems, it is crucial to examine the specific cyber threats that target these systems. Table 5 presents a chronological overview of both real-world attacks and emerging threats identified in the literature on AID systems.
As illustrated in the taxonomy (Figure 2), malicious attacks on AID systems can be broadly categorized into cyber-physical attacks and process-aware attacks, based on whether they exploit knowledge of the system’s dynamic control processes [52]. Cyber-physical attacks, which are common across various IoMT devices, compromise system functionality by targeting software, hardware, or communication channels without requiring an in-depth understanding of the underlying control mechanisms. In contrast, process-aware attacks specifically target semi- or closed-loop medical systems such as AID. By making subtle manipulations within the operational bounds of the system, these attacks can significantly impact patient safety [40]. Since cyber-physical attacks align with broader IoMT security concerns, threats and potential countermeasures have been detailed in existing surveys [29][30][31], this section focuses on analysing attack vectors and technical details that are relevant to AID systems.
To systematically analyse these attack vectors, we categorize them into three key areas based on their impacts on the fundamental principles of cybersecurity: confidentiality, integrity, and availability. Table 6 presents an overview of these attack vectors, along with brief descriptions and feasible countermeasures.
Attack Name Ref. Type Cause of Failure Impact Adversarial Capabilities RL-based Control Model Attack Chang et al. [53] (10/2024) Literature (Experimentally Validated) The control system is vulnerable to anomalous inputs. A maximum CGM reading of 400mg/dL caused high rates (over 40%) of hypoglycaemic failure within a few hours; a very slight perturbation (±1mg/dL) can lead to disproportionate patient harm over time High - Read access to the weights of the trained model or estimated weights AI/ML model attacks Elnawawy et al. [54] (01/2024) Literature (Experimentally Validated) Introducing adversarial CGM data points during inference through a known vulnerability in the Bluetooth communication channel Wrong diagnosis or treatment Medium - Can exploit the Bluetooth communication stack to intercept and manipulate the CGM measurements Unauthorized Remote Control Medtronic [10] (01/2023) Real-world (Recall) The settings and insulin delivery of the pump could be potentially controlled through a wireless connection nearby Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Process-Aware Attacks Stergiopoulos et al. [40] (11/2023) Literature (Experimentally Validated) Utilising knowledge of the dynamic process to alter sensory data or the logic inside the controller in an intended but subtle way Overdosing or Underdosing High - Prior knowledge of the components’ implementation mechanism Replay Attacks Walenda et al. [55] (11/2023) Literature (Experimentally Validated) Sending repeating insulin dosages to the insulin pump through a vulnerability in the wireless protocol Replay attacks have the greatest effect when the injected insulin doses are high Low – Intercept the wireless connection and report past packets DoS Attacks Walenda et al. [55] (11/2023) Literature (Experimentally Validated) Overload the system to stop insulin delivery for extended periods DoS attacks are most effective at low, increasing doses, causing longer out-of-range times and greater spikes and dips than replay attacks. Low - Jam the communication channel Personalized insulin dose manipulation attack Levy-Loboda et al. [11] (08/2022) Literature (Theoretical) Craft attacks that only threat target patients and are capable of evading extreme overdose/underdose detection Overdosing or Underdosing High – Prior knowledge and manipulation of the insulin dosage Bias Injection Attacks Emre Tosunal [50] (06/2022) Literature (Theoretical) A constant bias is injected into the compromised sensor readings Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Ripple False Data Injection Attacks Kashyap et al. [56] (05/2021) Literature (Experimentally Validated) Using minimal input perturbations to stealthily change the DNN-predicted insulin dosage More allowed perturbations make attacks easier to find for targeted attacks, with discovery time as short as 60ms. Random attacks can easily induce 10–20 mg/dL output changes across scenarios. High - Read access to the weights of the trained model or estimated weights Command Manipulation Attack Khan et al. [41] (12/2020) Literature (Theoretical) Change the instructions or data that implement the algorithmic steps Overdosing or Underdosing High - Access to the relevant resources; modify a command (i.e., instructions that implement the algorithmic steps) Unauthorized Remote Control Cooke et al. [13] (04/2020) Literature (Experimentally Validated) An adversary can remotely control insulin delivery and access unencrypted medical data Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Replay Attacks Medtronic [49] (08/2018) Real-world (Notification) Involuntary bolus insulin could be delivered to the pump user by copying the wireless radio frequency (RF) signals from remote controllers in proximity Overdosing Low – Intercept the wireless connection and report past packets CGM Attacks Reverberi and Oswald [57] (08/2017) Literature (Experimentally Validated) The Dexcom G4 protocol does not employ cryptographic algorithms Enables the tracking of a user, the jamming of the communication channels and the forging of incorrect sensor readings Low to medium - software-defined radio and low-cost RF chipsets, reverse engineering Saturation-based Sensor Spoofing Attacks Park et al. [58] (08/2016) Literature (Experimentally Validated) Make a sensor ignore legitimate inputs by using an additional infrared source to saturate the sensors Inject up to 3.33 times the intended amount of fluid or 0.65 times for a 10-minute period Low – Saturating sensor from an external source Software radio-based attacks Marin et al [48] (03/2016) Literature (Experimentally Validated) Fully reverse-engineered the wireless communication protocol between all the peripherals of the insulin pump system Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
Active attacks: Medium - Reverse engineering
Software radio-based attacks Li et al [12] (07/2011) Literature (Experimentally Validated) No standard cryptographic mechanisms; the wireless protocol can be fully reverse-engineered, enabling eavesdropping, replay, and command manipulation attacks. Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
Active attacks: Medium - Reverse engineering
Table 5: Documented attacks on AID systems.
Security Aspect Attack Vector Description Countermeasures Confiden-tiality Eavesdropping
[12][57]
Theft of information transmitted through a network Cryptography;
Proximity communication channel;
Secure protocols;
Authentication
Do-it-yourself hacking[16] Patients extract data not automatically provided by products’ software Authentication Integrity Bias injection attack[39][50] A constant bias was injected into the sensor readings Intrusion detection systems;
Redundant hardware or software components;
Infusion patterns mining;
Control strategies assess
False data injection attack[44][10] Change transmitted data with random or synthesized data Replay attacks[6] Intercept and replay legitimate data AI/ML model attacks[54] Deviate the ML model’s prediction during training or inference Computational code attacks[41] Compromise the command computational code Formal specification Malicious pump driver [59] Change pump I/O operations Avail-ability Denial-of-Service (DoS) /Distributed DoS [59] Flooding the target with traffic results in the inability to respond timely Authentication;
Secure protocols;
Formal specification;
Redundant systems
Ransomware Attacks [60] The malware encrypts the system’s data or locks users out, demanding a ransom to restore access Jamming Attacks [57] Disrupting the wireless communication channels that the APS uses to operate Firmware/ Software Corruption [60] Corrupting the firmware or software that the APS relies on to function Routing protocol information attacks [61] Attacks on routing protocols to disrupt or manipulate the flow of information
Table 6: Attack vectors and common countermeasures in AID systems.
4.1 Threats to Confidentiality
Ensuring the confidentiality of diabetes devices aims to guarantee that sensitive information is accessible only to authorized users or systems. Threats to this security aspect can originate from unauthorized individuals or even from patients themselves.
4.1.1 Eavesdropping
This refers to the unauthorized interception of wireless communication which can expose (1) personal patient information related to therapy and physical condition, (2) device information, and (3) the device’s personal identification number (PIN). Laboratory experiments have shown that data transmitted by a commercial insulin delivery system can be intercepted using off-the-shelf Universal Software Radio Peripheral (USRP) and free software [12]. By capturing its identification number, attackers could then gain full control of the pump. At a 2012 security conference, an engineer showed a fatal dose of insulin could be delivered from 300 feet away [62]. These attacks exploit wireless communication vulnerabilities, such as wide transmission range and inadequate authentication and encryption. To mitigate these risks, countermeasures proposed in the literature mainly focus on (1) enhancing data protection through secure authentication, cryptography, and communication protocol, and (2) limiting transmission range by employing proximity communication channels.
4.1.2 Do-it-yourself (DIY) hacking
This refers to unauthorized access or modification of commercial diabetes devices by patients themselves driven by the desire for better data visualization and more advanced control with shorter official censor and approval time [16]. For instance, Nightscout is an open-source project that helps patients hack their Dexcom CGMs for remote glucose level monitoring on mobile devices. Projects such as DIYPS and #OpenAID systems enable patients to control insulin delivery. While these initiatives empower patients, they introduce risks due to vulnerabilities in unauthorized software and users’ potential lack of expertise. Compared to FDA-regulated medical software, these systems lack rigorous testing and responsible parties to support them in case of problems. To balance the need for patients to have greater control with secure AID systems against malicious exploitation, it is crucial to implement robust authentication strategies.
4.2 Threats to Integrity
The integrity of AID systems ensures that its data, algorithms, and operations remain accurate and consistent. Integrity can be compromised through two key aspects: data manipulation and algorithm manipulation.
4.2.1 Manipulation of Data
Attackers can manipulate wirelessly transmitted data, such as CGM readings or controller commands, to mislead the control algorithm or insulin pumps to issue incorrect insulin delivery. Data manipulation can occur through the following attack vectors:
  • 1.
    Replay Attacks: These attacks involve intercepting and replaying legitimate data packets to manipulate insulin delivery [6]. This type of attack is relatively simple to launch as it does not require device access and message decryption. Once the attacker gains access to the device (e.g., access the PIN), they can craft forged data packets to be accepted by the device [12], enabling bias injection and false data injection attacks.
  • 1.
    Bias Injection Attacks: In this attack, the attacker skews sensor readings by injecting a bias [39]. By making subtle, consistent changes, the attacker can cause the system to under or over-deliver insulin without needing extensive knowledge of the patient model or system operations. They can be difficult to detect due to their subtlety and correlation with the original data [63].
  • 1.
    False Data Injection Attacks (FDIA): FDIA could be categorized into random and targeted types. Random FDIA introduces arbitrary interference in the input data. Experiments in [56] showed that such attacks can cause outputs of ML-based glucose prediction models to change 10-20 mg/dL. This type is easier to detect through sequential change detection systems. In contrast, targeted FDIA synthesize precise input values to evade alarm and cause targeted harm to patients. Launching such attacks requires knowledge of the model’s architecture and parameters, but this information is often detailed in documentation. Kashyap et al. [56] formulated this attack as an optimization problem, demonstrating that an effective targeted FDIA with minimal, hard-to-detect perturbations can be computed within 60 milliseconds.
Defence strategies for these data manipulation attacks typically include developing intrusion detection systems to detect abnormal patterns in sensor data. Another defensive measure involves using redundant sensors or software monitors to cross-check data sources for inconsistencies.
4.2.2 Manipulation of Algorithms
Algorithm manipulation involves tampering with the computational logic or parameters setting of AID systems, targeting either control algorithms or physiologic models used for patient state prediction [41]. These include both traditional rule-based algorithms and machine learning-based models.
  • 1.
    Computational code attacks: These attacks compromise the command computational code in rule-based algorithms [64]. For example, Aliabadi et al. [65] identified that the synchronization between CGM and controller could be manipulated by altering the timing function (get-data-timer()). This modification allows attackers to delay or drop critical messages. Additional threats include altering essential settings, such as disabling pump alarms, modifying the maximum dosage limits, or enabling insulin overdoses [40][66][59].
  • 2.
    ML model attacks: ML models are increasingly used in AID systems to enhance performance [67]. For instance, the FDA-approved One Drop Blood Glucose Monitoring System employs a multi-layer perceptron (MLP) model for diabetes management. Deep reinforcement learning (RL) approaches have been explored for use in AID systems for BG control [68]. However, due to their black-box and data-dependent nature, understanding and monitoring the reasoning behind these models is challenging. Attackers may influence the model’s behaviour during training or inference phases or reverse-engineer the model to extract confidential information [54]. Moreover, ML-based algorithms can be severely interfered by a very slight input data perturbation that is hard to detect. In [56], minimal interference was calculated to misguide the ML model into predicting a patient’s glucose level as hyperglycaemic, leading the controller to administer an excessive dose of insulin. Experiments in [69] show that a small perturbation around ±1mg/dL on BG reading can lead to severe patient harm over time in an RL controller model.
To defend against algorithm integrity threats, formal verification techniques are often employed to assess the operations of rule-based algorithms. For ML-enabled systems, which pose greater challenges for formal verification, patient-specific insulin injection patterns have been modelled to detect abnormalities and inconsistencies.
4.3 Threats to Availability
Availability in the context of AID systems refers to ensuring that the system and its components remain reliably functional for data transmission, processing, and regular operations. Threats to availability potentially lead to failures in glucose monitoring or delays in insulin delivery. Key threats include:
  • 1.
    Denial-of-Service (DoS) and Distributed DoS (DDoS): In these attacks, an adversary floods the system with excessive traffic, overwhelming AID systems and preventing legitimate requests from being processed promptly. This could lead to delays or complete failure in delivering insulin.
  • 2.
    Jamming Attacks: By interfering with the wireless communication frequency, attackers can interrupt data flow, causing the system to fail to deliver insulin as required.
  • 3.
    Firmware/Software Corruption: Malicious firmware or software corruption can cause AID systems devices to malfunction or stop functioning entirely. Attacks exploiting software vulnerabilities or delivering harmful updates can prevent systems from operating as intended, leading to a cessation of insulin administration.
  • 4.
    Routing Protocol Information Attacks: Manipulating data transmission pathways in networked AID systems, attackers can disrupt or delay communication between key system components, compromising timely insulin delivery and glucose monitoring.
  • 5.
    Ransomware Attacks: Ransomware can encrypt data and lock critical components of AID systems, demanding ransom payments to restore access. This can prevent patients or healthcare providers from viewing essential data, such as glucose levels or delivery history, rendering the system inoperable until resolved.
Mitigation strategies for these threats include network traffic monitoring, secure communication protocols, authentication mechanisms, and the application of formal verification methods.
5 Defence Strategies for AID systems
Addressing the wide range of potential cyberattacks on AID systems requires a comprehensive defence strategy that spans multiple system layers. Defence mechanisms in AID systems can be grouped into three primary areas: protected communication channels, intrusion detection, and control activities monitoring. Besides these internal defences, external strategies—such as the use of redundant hardware or software components—are employed to detect abnormal activities independent of AID devices. Table 7 provides an overview of the defence types and corresponding methods proposed in past research for protecting AID systems.
Defence type Description Methods Protected communi-cation Authentication Verify the identity of users or devices Fingerprint recognition[70][71];
Voiceprint analysis[72];
Physical Unclonable Functions[73][74];
Blockchain[73][75]
Cryptography Transform information into an unreadable format for unauthorized users Rolling-code encryption[12];
AES-based encryption[76][57];
Asymmetric encryption [70];
Homomorphic encryption [77]
Secure communication protocols Define rules for secure data transmission RF-Harvested energy-based protocol[60];
Emergencies-resilient protocol[37];
Improved the Dexcom G4 protocol[57];
Mutual authentication protocol[78]
Proximity communication channels Require close physical proximity between the communicating devices Body-coupled communication[12];
Visible light communication[79]
Intrusion detection systems Signature-based Compare the real-time behaviour of the system against known security attacks Define computational, data integrity, and communication attacks[41];
Define abnormal signal’s physical and content features[5];
Define replay and DoS attacks[55];
Infusion site failures [80]; Pressure-induced sensor attenuations [81]
Specification-based Define specifications of AID systems’ normal operations and deem any violation as an intrusion Formal specification [82] with
personalized thresholds[83][59];
statistical analysis[84];
declarative property [41];
outlier detection[34][85];
specification mining[86];
external and internal state[87][88];
insulin predictive model[59]
Anomaly-based Identify data patterns that have different data characteristics from normal instances Model-based methods:
Physiological [89][90][50][91][92][93][94];
Data-driven [95][96][97][98][99][75][100] [33][101][102][103];
Model-free methods:
Supervised [104][105];
Unsupervised [43][106][107][108][109];
Hybrid methods [110][34][85][59]
Control strategies assessment Compare the dosage with the identified patient’s insulin infusion pattern Rule-based [41];
Data-driven[66][11];
Hybrid[83]
Redundant hardware or software components Use of multiple, often independent, components that perform the same function Hardware:
ECG monitor[88]; Universal Software Radio Peripheral[5]; Separate computation core[60]
Software:
Replay attack detection filters[6]; Auxiliary safety layer for controllers[111]
Table 7: Defence strategies proposed for AID systems.
5.1 Protected Communication
Cyberattacks often exploit vulnerabilities in wireless communication channels. A variety of defence mechanisms have been proposed to enhance the security of information transmission within AID systems, including advanced authentication methods, encryption techniques, secure communication protocols, and specialized communication channels.
5.1.1 Authentication Methods
Reliable authentication is crucial for ensuring secure communication within AID systems and for only allowing remote control with trusted devices. Several authentication methods have been explored to meet the high-security and low computational demands of AID systems.
Biometric authentication methods like fingerprint recognition [70][71] and voiceprint analysis [72] are user-friendly and offer strong uniqueness but introduce concerns about sensitive biometric data storage, which could be spoofed through advanced attacks. To mitigate these risks, hardware-based security methods such as Physically Unclonable Functions (PUFs) exploit inherent and unique physical variations in manufacturing processes, offering low computational overhead and an unclonable security mechanism without the need for key storage [73][74]. Additionally, blockchain technology has been explored to protect patient data by decentralizing data storage in central cloud servers [73]. Rahmadika et al. [75] further incorporated blockchain with federated learning for sharing local parameters for the misbehaviour detection model in AID systems without compromising patient privacy.
5.1.2 Encryption Techniques
Following successful authentication, establishing encrypted communication between AID systems components is vital for maintaining data confidentiality and integrity. A key challenge is the extreme energy and computational constraints. High-complexity encryption algorithms can introduce excessive computational overhead, reducing transmission frequency and accelerating battery depletion—both of which could hinder the system’s ability to respond promptly to life-critical scenarios. Another challenge is key distribution. While private key encryption enhances security, it may prevent medical professionals from accessing patient data in emergencies [5].
To address these challenges, several studies have proposed encryption methods specifically tailored for AID systems. Table 8 summarizes these approaches, detailing their methodologies, key sizes, security strengths, and limitations.
Encryption Method Description Key Sizes Security Strength Limitations Rolling code en/decoder [12] The remote controller encrypts an incrementing sequence counter as a shared key which is decrypted by the insulin pump to verify the counter within an acceptable range 32–128 bits (Keeloq: 64-bit) Provide basic security by preventing straightly intercepting the device series number and replay attacks Not offer strong authentication. Not clarifying the management of the shared key. AES-128 optimized for energy consumption [76] The remote control and the insulin pump share two independent symmetric cryptographic keys for encryption and authentication. Message size and transmission frequency are reduced to lower communication costs. 128 bits Strong security and relatively lower computational cost Vulnerable if public keys are intercepted ForkAE [13] Simultaneously performs encryption and authentication by creating two copies of ciphertext for the plaintext but adds an extra constant to one of the copies and encodes both blocks using the same encryption algorithm and key 128 bits Strong security and lower computational cost than AES-128 Vulnerable if the fork pattern is predictable or public keys are intercepted RSA encryption using SHA generated 1024 key [70] An asymmetric cryptographic pattern is applied between a smartphone (communicate with sensors and actuators) and a control laptop 1024 bits Moderate security. Include data signing with the private key in every packet in case public keys are intercepted. Higher computational cost. Prone to factorization and quantum attacks A Fully Homomorphic Encryption scheme (CKKS) [77] A PID controller was employed to use an encrypted glucose signal as an input to produce encrypted control signals Variable, typically in the range of hundreds of KB to MB Very strong security and preserved privacy. Computational cost is very high. Lack of division and comparison operations.
Table 8: Proposed cryptography-based approaches in AID systems.
Li et al. [12] first demonstrated the vulnerability of wireless communication between glucose monitors and insulin delivery systems due to the lack of standard cryptography, they proposed using rolling code-based encryption—similar to that used in garage door systems—to prevent eavesdropping and replay attacks. However, the security of rolling code methods does not offer strong authentication and is highly dependent on encryption and decryption algorithms. Some methods, such as Keeloq, have been successfully compromised [112]. To enhance security, Marin et al. [76] employed Advanced Encryption Standard (AES), a more robust symmetric encryption algorithm. Additionally, they reduced message size and transmission frequency to optimize computational efficiency. Their experiments also revealed that authentication for messages consumed more energy than encryption due to the need to append a message authentication code (MAC), which inevitably increased message length. A more lightweight solution proposed by ForkAE [100] combined encryption and authentication processes to reduce computational overhead.
Both AES and ForkAE rely on symmetric encryption, where a public key is used for both encryption and decryption. However, this approach carries the risk of key interception. To mitigate this problem, asymmetric encryption was introduced in [70], using a private key in every packet. While this approach enhances security, it requires larger key sizes, resulting in higher computational costs and vulnerability to factorization attacks. To further protect patient privacy during computation, Weng et al. [77] proposed using homomorphic encryption within a PID-based controller. This approach enables secure data processing without key sharing, as it allows computations to be performed on encrypted data. However, the CKKS homomorphic encryption method hired in their study does not support division and comparison operations, potentially leading to imprecise computational results. Moreover, its high computational demands significantly limit its application to more complex controller algorithms.
5.1.3 Secure Communication Protocols
A communication protocol defines the rules and standards for data exchange between devices, including authentication and encryption methods, to ensure secure and efficient transmission. Traditional high-security communication protocols often fall short in AID systems due to their substantial computational and energy demands. In contrast, low-power communication protocols like Bluetooth Low Energy (BLE), commonly used in IoT devices, often lack security robustness.
When applied in AID systems, the absence of authentication mechanisms in BLE to verify message integrity makes it vulnerable to man-in-the-middle and replay attacks. To address this concern in AID systems, Kwon et al., [78] proposed a protocol that introduces mutual authentication with 35.7% less computational overhead compared to BLE. Except for authentification, the original protocol of a CGM (Dexcom G4) has been demonstrated to lack encryption [57], leaving it susceptible to eavesdropping and jamming attacks. To mitigate these threats, Reverberi and Oswald enhanced the CGM (Dexcom G4) protocol by incorporating AES-128 encryption and authentication [57]. Furthermore, the use of random keys and frequencies was introduced to protect against jamming attacks. While this approach mitigates simple narrow-band jamming, it remains susceptible to high-power wide-band jammers and more advanced jamming techniques. Strydis et al. [60] introduced an energy-efficient protocol that runs on a separate core powered by RF-harvested energy until it performs external reader authentication. This approach reduces the energy burden and enhances the system’s resilience to battery DoS attacks.
5.1.4 Proximity Communication Channels
Besides protecting the data transmission process, limiting transmission range through proximity-based communication can reduce the risk of data interception and manipulation. Li et al. [12] proposed hiring body-coupled communication (BCC) that uses the human body as the wireless communication medium, which reduces wireless interception risks and power consumption. However, BCC can be affected by variations in body conductivity and movement. An alternative is visual light communication [79], which transmits data over short distances using visible light. This approach makes interception difficult and prevents remote attacks from hidden locations but is sensitive to obstacles and ambient light, limiting its use to controlled environments like hospitals.
5.2 Intrusion Detection Systems
Monitoring and analysing system activity and states is vital to detect any intrusion and abnormal behaviours, ensuring the proper functioning of AID systems. Intrusion Detection Systems (IDS) have emerged as a key defence mechanism to counter threats. These systems are classified into three primary types: signature-based, specification-based, and anomaly-based approaches. Table LABEL:table4-3 provides an overview of studies that have developed IDS solutions specifically for AID systems or have used AID systems as a case study.
Table 9: Intrusion detection systems developed for AID systems.
Reference Category Detection method Threats Evaluated Dataset Performance metric
Zhang et al. (2013) [5] Signature-based Monitor all the RF wireless communication. Identify and jam potentially malicious transactions Unauthorized remote control Simulated with glucose sensor and insulin delivery system Detection rate: 100%;FNR: 0% Detection time: about 2.5 ms
Khan et al. (2020) [41] Signature-based Specify each attack as a one big step relation between the initial and final state Controller command manipulation attacks Software simulation on a 2.6-GHz Intel Core i7 processor CPU Time: 7.50 ± 0.3 * 10-4 s; Run time: 7.60 ± 0.3* 10-4 s
Walenda et al. (2023) [55] Signature-based For replay attacks, detect consecutive repeated values. For DoS attacks, detect consecutive 0. Replay attacks; DoS attacks T1D patient simulator - GlucoSym Simulator Qualitative illustration of detection
Howsmon et al. (2018) [80] Specification-based Patient-specific metrics are calculated and compared to sliding window averages. Alarm when thresholds are exceeded. Infusion site failure 19 participants, 2 weeks Sensitivity: 88.0% (n = 25) FPR: 0.22 per day
Alshalalfah et al. (2019) [84] Specification-based Combine formal methods with statistical analysis to evaluate the safety of control algorithms Replay attacks T1D patient simulator - UVA/PADOVA 2013 (5 patients, 24 hours) Qualitative evaluation
Aliabadi et al. (2021) [86] Specification-based Dynamically mining specifications based on their contribution to detecting every arbitrary attack CGM spoofing attack; Stop basal injection; Resume basal injection OpenAPS simulator FPR: 13.5%; FNR: 2%
Astillo et al. (2021) [34] Specification-based Incorporated specification with an outlier detection method Manipulate random points of the glucose data within 1% to 10% deviation UVa/Padova simulator (2008) + Raspberry-Pis AUROC: Attack rate 50%-90%: 99.94% ± 0.04%; Attack rate 10%-30%: 54.09%± 2.5%
Astillo et al. (2021) [85] Specification and anomaly-based Extend SMDAps with timing-based malicious transmission detection and trustworthiness evaluation Manipulate random points of the glucose data within 0% to 50% deviation UvA/Padova simulator (2008) + Raspberry-Pis (2 virtual patients, 24 hours) Accuracy under the hidden attack mode: 99.1% ± 0.7%
Prematilake et al. (2021) [87] Specification-based Combined state transition rules, I/O access rules, and physiological rules to monitor both extrinsic state and internal state Recurring bolus dose; Bolus unmatched with meal taking; Buffer overflow Simulated with a safety-enhanced controller board and glucose-insulin model Delay time: 253 ms; Energy overhead: 3% (a 2500 mAh AA-sized battery)
Zhou et al. (2021) [83] Specification-based Formal specified unsafe system context (physiological) and personalized thresholds using reinforcement learning Availability attack; DoS attack; Integrity attack/memory fault Simulated with Glucosym and UVA-Padova Simulator 2013 F1-Score: 0.975 ± 0.005
Venugopalan et al. (2024) [59] Specification and anomaly-based Combined formal methods (prove correct the most critical parts) and dictated safe bounds of insulin amounts to enable DL/ML to control the insulin pump safely. Malicious ML; Malicious pump drivers Real-world (1 individual, 6 months); UvA/Padova simulator (2008) (21 virtual humans) Time In Range (TIR): Adults:92.88% Adolescents: 82.38%; Children: 85.42%
Vega-Hernandez et al. (2009) [89] Anomaly-based (physiological model) Incorporated the effect of the interstitial glucose measurements and the meal intake dynamics into the glucose-insulin metabolism model Over- and sub-dosing MATLAB SIMULINK simulation, 40 days with a normal patient routine Qualitative evaluation
Facchinetti et al. (2013) [95] Anomaly-based (Data-driven model) An individualized state-space model and a Kalman predictor 1) spike on CGM 2) a loss of sensitivity of the CGM sensor 3) stop insulin delivery UVA/Padova 2008, 100 simulated patients; 3 real T1DM subjects CGM failure: FNs: ¡ 5%, FPs: ¡ 10%, CSII pump failure: Detection rate: 75% within average 60 min, FPs: 10%
Favero et al. (2014) [96] Anomaly-based (Data-driven model) Extend [95] for whole-day use by taking meals into account 1) spike on CGM 2) consecutive large errors 3) meal*(1+e%) 4) bias meal bolus 5) basal faults UVA/Padova 2008, 100 simulated patients Sensitivity: CGM faults:   85% Meal faults:   90% Bolus faults:   90% Basal failure: 50%
Meneghetti et al. (2020) [97] Anomaly-based (Data-driven model) Distinguish insulin pump faults and missed meal announcements by introducing 2 sets of parameters of model Night-time pump occlusion ; missed meal announcement ; meal-time pump occlusion UVA/Padova Type 1 Diabetic Simulator 2013 Pump faults: sensitivity:  81.3% FPR:  0.15/day missed meals: sensitivity:  86.8% FPR:  0.15/day
Meneghetti et al. (2023) [98] Anomaly-based (Data-driven model) Extend [97] with analyse various statistical properties Insulin suspension; CGM missed samples and pressure-induced sensitivity losses; missed meal announcements UVA/Padova Type 1 Diabetic Simulator 2013 Recall: 91.3% FP: 1 every 10 days
Tosun et al. (2022) [50] Anomaly-based (physiological model) Kalman filter and a $\chi^{2}$ test. Identify meal disturbance by an online meal estimator and a time-varying threshold Bias injection attack on glucose sensor Medtronic Virtual Patient (MVP) model (1 virtual subject) + PID controller Qualitative evaluation
Tosun and Teixeira (2023) [91] Anomaly-based (physiological model) Kalman filtering and sequential change detection (detecting an abrupt change) deterministic sensor deception attacks (a slowly increasing bias injection attack Bequette model (glucose) Qualitative evaluation (the $\chi^{2}$ test maximizes the best-case detection delay, while the CUSUM test minimizes the worst-case delay)
Tosun et al. (2024) [92] Anomaly-based (physiological model) The prediction error of a time-varying Kalman filter is statistically evaluated to detect anomalies under meal disturbances with known time Sensor bias injection attack Medtronic Virtual Patient (MVP) model (1 virtual subject) + PID controller Qualitative evaluation
Turksoy et al. (2016) [102] Anomaly-based (Data-driven model) Non-linear first principle dynamic model with an unscented Kalmen filter. PCA and K-means clustering hired for fault detection Faulty CGM measurements Data from 51 subjects Sensitivity: 84.2 %; Average detection time: 2.8 min
Olney et al. (2022) [99] Anomaly-based (Data-driven model) Non-linear autoregressive neural networks Incorrect sampling of the target biosignal (sensor data) D1NAMO CGM dataset (9 human subjects) + FPGA Average RMSE: Software: 0.428 Hardware: 0.455
Maity et al. (2023) [100] Anomaly-based (Data-driven model) Learn surrogate model for safety-certified actions using RNN, then periodically relearn model coefficients to check deviation Insulin cartridge problem UVA/Padova 2018 Difference in model coefficient values: $>$ 0.27
Rahmadika et al. (2023) [75] Anomaly-based (Data-driven model) BiLSTM and augments the security through integrating blockchain technology Arbitrarily changing a feature value (increase or decrease 10% - 50%) UVA/Padova 2008 Average RMSE: 1.3482 (5 timesteps), 1.9959 (7 timesteps), 2.6099 (9 timesteps)
Zhou et al. (2022) [101] Anomaly-based (Data-driven model) Extract the specification of unsafe control actions and add them as regularization terms in a semantic loss function Designed small but malicious input perturbations for both sensor and commands Glucosym simulator; UVA-Padova Simulator 2013 Robustness error: Reduce on average up to 54.2%
Meneghetti et al. (2018) [109] Anomaly-based (Model free) Local outlier factor and connectivity-based outlier factor AD methods insulin pump failure(suspensions occurring at night for 6 hours) UVA/Padova Type 1 Diabetic Simulator 2013 Precision:  75%; Recall  60%
Meneghetti et al. (2019) [106] Anomaly-based (Model free) Unsupervised anomaly detection with optimum feature set selection Insulin pump faults Padova/UVA T1D simulator 2018, 100 adults, 30 days Histogram-based outline score: Sensitivity: 0.87 FPR: 0.08/day Isolation forest: Sensitivity: 0.85 FPR: 0.06/day
Meneghetti et al. (2022) [43] Anomaly-based (Model free) Unsupervised anomaly detection Infusion site failure Clinical dataset (N=20) Sensitivity: 0.75 (15 out of 20); FPR: 0.08/day
Meneghetti et al. (2020) [57] Anomaly-based (Model free) Unsupervised multi-dimensional data-driven anomaly detection. Cut several start and end points to enlarge the difference between normal and abnormality Suspension of insulin delivery UVA/Padova T1D Simulator 2013; 7 patients, 1 month in free-living conditions Recall:  80% Precision: $>$90%
Herrero et al. (2022) [108] Anomaly-based (Model-free) Detect whether CGM data belong to the individual using SVM binary classifier with 12 standard glycemic metrics at 24h, day, night, breakfast, lunch, and dinner Reidentification REPLACE-BG (NCT02258373), 226 adult, 6 month Accuracy: 86.8% ± 0.16; F1-score: 0.86 ± 0.16 (window length: 15 days)
Astillo et al. (2022) [71] Anomaly-based (Data-driven and Model-free) Deep learning-based anomaly detection system with federated learning False sensor data UVa/Padova simulator (2008) Accuracy: 99.17%; F1-score: 99.16%
Judith et al. (2023) [105] Anomaly-based (Model free) PCA for feature reduction and multi-layer perceptron (MLP) for classification Man-in-the-middle attacks WUSTL-EHMS-2020 dataset Sensitivity: 95.4% Specificity: 100%
5.2.1 Signature-based IDS
Signature-based IDS detect malicious activities by comparing real-time system behaviour against a database of known attack patterns. For example, Zhang et al. [5] detect malicious data transactions in AID systems by defining abnormal physical signal characteristics and behaviours of the underlying content. The patterns of replay and denial-of-service (DoS) were defined in [55] and computational, data integrity, and communication attacks were defined in [41]. For unintended failure detection, Howsmon et al.[80] hired patient-specific metrics to monitor infusion site failures and Baysal et al. [81] defined sets of rules for detecting pressure-induced sensor attenuations.
One key advantage of signature-based IDS is its lightweight computational overhead, making it well-suited for resource-constrained medical devices. However, it is ineffective against zero-day (unknown) attacks. Besides, updating signature databases in AID systems presents a challenge due to the stringent regulatory approval processes for medical devices. Despite these limitations, signature-based IDS can serve as a rapid first-line defence by filtering known threats efficiently and combining with other IDS capable of unknown attacks to offer more adaptive defences.
5.2.2 Specification-based IDS
Specification-based IDS identify malicious activities by comparing system behaviour to predefined rules or models that outline the expected operations of the device. Several studies have been conducted to improve formal specifications for AID systems, aiming to enhance precision and efficiency.
To mitigate the high resource demand of specification-based IDS for real-time performance requirements of AID systems control, Chen et al. [82] proposed a simplified multi-basal insulin infusion control model for verification, which addresses time-triggered jumps and merges flow pipes over the same time interval. Alshalalfah et al. [84] integrated statistical analysis with formal methods, employing statistical model checking to minimize exhaustive state exploration. Subsequently, they proposed a dynamic specification mining approach using Bayesian networks to refine rules, reducing computational load while maintaining detection accuracy [86]. Khan et al.[41] introduced a declarative property-based approach, specifying only the initial and final states rather than all intermediate states in a runtime monitor.
Rigid thresholds present another issue, as they may not adapt well to individual patient variability. Zhou et al. [83] tackled this by personalizing thresholds through adversarial training on fault-injected data, enabling more precise monitoring. Beyond internal system states, unsafe AID systems operations can manifest as external abnormalities, such as repeated bolus requests or mismatches between insulin delivery and meal intake. Prematilake et al. [87] developed a rule-checking mechanism that monitors both extrinsic and internal states.
Ideal specification-based IDS are effective in defending against unknown attacks. However, constructing comprehensive rule sets to capture the complex and dynamic interactions within closed-loop insulin delivery systems and limiting computational overhead remain significant challenges in practice.
5.2.3 Anomaly-based IDS
Anomaly-based IDS identify deviations from expected data patterns to detect abnormal activities. These approaches are generally categorized into model-based and model-free methods.
Model-based approaches consider a T1D patient as a dynamic system, where the glucose concentration $g(t)$ is influenced by the insulin injection $i(t)$ and process disturbances $m(t)$ (e.g., meals and other patient activities). Intrusions are detected by comparing model prediction with observations. Based on the type of model, these approaches are further split into physiological model-based and data-driven model-based methods.
  • 1.
    Physiological Model-based: These methods utilize mathematical models to represent T1D patient profiles, often integrating predictors such as Kalman filters for state estimation. Early work by Vega-Hernandez et al. [89] leveraged Hovorka’s nonlinear predictive model [113], while Avila and Martínez [94] used Lehmann and Deutsch model combined with a stochastic process to detect insulin pump issues.
    Several studies have focused on improving the prediction accuracy of these models. Tosun et al. [91] added sequential change detection mechanisms to prompt detection speed and proposed a KL-divergence bias-sensitive filter to improve the detection rate of constant bias injection attacks [93]. To distinguish unregistered meals that caused false alerts, they introduced an online meal estimator using $\chi^{2}$ test with adaptive thresholds [50], which was later enhanced by including the statistical evaluation of prediction error [92]. A comparison of three nonlinear predictive filters for online drift detection in BG data was presented in [90].
    The advantages of physiological model-based approaches include the use of expert knowledge, explainability, and a reduced requirement for extensive patient data. However, their accuracy is highly dependent on the quality of the underlying model.
  • 1.
    Data-driven Model-based: These approaches rely on learning glucose-insulin dynamics from collected data. Early methods primarily employed statistical or linear models to represent these dynamics. For example, multivariable statistical monitoring methods were used in [102] for detecting faults in glucose concentration values. Facchinetti et al. [95] proposed individualized linear models to improve glucose predictions at night, which Del Favero et al. [96] later expanded to cover day-long monitoring by incorporating meal effects. To distinguish between unannounced meal interference and pump malfunctions, they introduced two different sets of parameters in a personalized autoregressive moving average model [97]. This work was further enhanced by incorporating a comprehensive analysis of various statistical properties associated with the residuals [98].
    In recent years, machine learning (ML) and deep learning (DL) methods have been applied to model glucose-insulin dynamics. For instance, Olney et al. [99] employed nonlinear autoregressive neural networks to detect and correct errors in CGM data, while Mahmud et al. [33] individualized this model to enhance prediction performance. Rahmadika et al. [75] developed a BiLSTM-based two-tier system for anomaly classification based on predictive deviations. Beyond modelling glucose-insulin dynamics, Maity et al. [100] proposed to model the system’s operational characteristics and detect attacks by periodically relearning the model and comparing coefficients.
    Though high detection accuracy is achieved with these models, they often lack explainability and may not perform well on out-of-distribution data. To address these limitations, Venugopalan et al. [59] suggested complementing ML-based models with personalised physiological models to enhance robustness. Zhou et al. [101] incorporated unsafe action specifications into the training process using semantic loss regularization. Building on this approach, they later introduced a data-driven hazard mitigation method that uses the desired sequence of recovery states to generate a corresponding sequence of control actions [103].
A key challenge faced by model-based IDS is that their performance heavily depends on the accuracy of the patient model. However, the model highly varies between patients and fluctuates over time [83]. Fixed models that do not account for these dynamic changes will struggle to maintain consistent performance over time. This underscores the need for personalised models that can continuously adapt to patients’ evolving physiological state to ensure reliable and effective anomaly detection.
Model-free anomaly detection methods leverage data-driven classification techniques to identify abnormal patterns without requiring a patient-specific physiological model. These methods are divided into supervised and unsupervised approaches based on the need for labelled abnormal data.
  • 1.
    Supervised Learning: These methods train classifiers on labelled datasets to recognise faulty or attacked data. Rojas et al. [104] mapped labelled data onto a feature map and defined boundaries for the “faulty” class, flagging patient glucose trends that enter this region. Judith et al. [105] employed an MLP model for classifying biometric data and network flow metrics. Despite their effectiveness, supervised methods face challenges due to the rarity and difficulty of obtaining labelled faulty or attacked data in the real world.
  • 2.
    Unsupervised Learning: To address the lack of labelled data, Meneghetti et al. [109] used 2 unsupervised anomaly detection methods: Local Outlier Factor and Connectivity-based Outlier Factor. Then they evaluated several unsupervised algorithms with feature selection techniques on simulated AID systems malfunction data [106]. Later, they extended this work to real patient data, enhancing the differentiation between normal and abnormal by adjusting the data segmentation [107]. Their follow-up study used a larger clinical dataset and adapted features for cases without meal announcements [43]. Herrero et al. [108] demonstrated that training a support vector machine to classify CGM data streams could accurately identify individual patient data. This method potentially aids in the detection of compromised data by reidentification.
While model-free methods bypass the need for profiling a physiological model, they still face challenges like managing false positives, optimizing feature selection, and handling distribution shifts to maintain effective anomaly detection in AID systems.
5.2.4 Hybrid methods
Hybrid methods combine various IDS approaches to enhance detection accuracy and robustness. For instance, Astillo et al. [110] merged model-based and model-free approaches, using deep learning models to forecast glucose levels and classify anomalies. They also incorporated outlier detection into specification rules to capture deviations in glucose data [34][85]. Venugopalan et al. [59] combined formal verification of critical system components with machine learning for insulin prediction, applying patient-specific safety thresholds. These approaches leverage the precision of specification-based or model-based methods and the adaptability of machine learning for more robust AID systems security.
5.3 Control Strategies Assessment
To prevent abnormal insulin delivery resulting from compromised CGM data, controller algorithms, or insulin administration, various defence strategies have been developed to evaluate and verify insulin dose control mechanisms. These strategies can be categorized into rule-based, data-driven, and hybrid approaches.
5.3.1 Rule-based methods
Rule-based methods compare insulin doses against predefined thresholds or patterns for anomaly detection. Khan et al. [41] used static threshold ranges to detect overdoses or underdoses, which, while effective for extreme deviations, fail to capture sophisticated attacks that account for patient variability. To improve this, Zhou et al. [83] introduced personalized thresholds and specific abnormal behaviours to align control actions with patient needs. However, this method detects anomalies after the patient’s physiological response to attack becomes apparent, leaving less or no time to prevent harm, especially in the case of underdosing, where adverse effects manifest more subtly.
5.3.2 Data-driven methods
Data-driven methods detect anomalies by predicting insulin doses using historical data, aiming to enhance timeliness and accuracy. Hei et al. [66] employed supervised learning to model normal infusion dosages and rates against time. However, their approach assumes rigid patient routines, limiting applicability, especially in pediatric cases. Levy-Loboda et al. [11] exploited more data sources such as the logs of the insulin pump, CGM and relevant medical information. They simulated insulin dose manipulation attacks and mined temporal patterns to identify anomalies using Logistic Regression and Random Forest. However, their study focused only on nighttime basal injections, neglecting other activities like meals and bolus doses.
Comprehensive models integrating patient activities, meals, and physiological changes for accurate and robust profiling of insulin patterns remain challenging.
5.4 Redundant Hardware or Software Components
[…]
While these external defences provide added protection without affecting device performance, they are vulnerable to initial compromise and can be combined with internal security measures for comprehensive AID systems protection.
6 Experiment Settings and Available Resources
Evaluating security systems for AID systems requires datasets that capture both the effects of cyberattacks on device functionality and the resulting physiological changes in patients. However, collecting such data through experiments is unpractical due to safety concerns. Besides, clinical data involving suspicious behaviour linked to cyberattacks is rare and difficult to label. These limitations often necessitate the use of simulation tools for testing and validating defence mechanisms. This section reviews evaluation methods and key experiment resources, including hardware, simulators, and datasets used in AID systems security research. The details of key experiment resources are provided in 10, which includes descriptions, sources, and relevant applications.
Type Resource Name Description Focus area Availability Usage in Research Hardware simulation Commercial diabetes medical products Widely used hybrid closed-loop insulin delivery system or its components Device functionality off-the-shelf Identify cyber vulnerabilities in APS [12][48];
Evaluate cryptography and authentication [70]
Universal Software Radio Peripheral (USRP) A software radio board that can intercept radio communications and generate wireless signals Wireless communication off-the-shelf Hijack APS [12][48];
Monitor wireless transactions[5]
Biostator II[114] An ultra-low-power ASIC implementation of an AP controller Implant functionality Open source Evaluate the energy efficiency of security communication protocol[60] Microcontroller board Simulate controller algorithms on board (e.g., Arduino Uno-R3[79][115][116]; EFM32WG[87], Arduino Nano 33 kit , nRF52840 development kit, and Galaxy S20 with Android 13[37]; ESP8266 development kit [117] Integrated function off-the-shelf Incorporated visible light access control channel [79];
Implement rule check mechanisms [87]
T1D patient simulator UVA/Padova 2008 [118] An FDA-approved nonlinear simulator captures interpatient variability [119] In silico trials Open source Simulate attacked monitoring data and evaluate intrusion detection systems[95][96] [34][110][59] UVA/Padova 2013 [120] Changes patient’s physiology over time In silico trials Licensed Evaluated control algorithms[84];
Unsupervised anomaly detection [107];
Homomorphic encryption[77];
Supervised anomaly detection[117]
UVA/Padova 2018 [121] Introduce intraday variability of insulin sensitivity, time-varying distributions of the patients’ therapy parameters, and a “dawn” phenomenon model In silico trials Not public open Feature selection and evaluate unsupervised anomaly detection algorithms[122]; Early detection of unknown-unknowns attacks [100] Glucosym [123] Online simulator that allows different control algorithms to be integrated to test their efficiency with 10 real-world patient profiles In silico trials Open source Simulate attack and evaluate defence methods [55] Pharmaco-kinetic model Hybrid physiologically-based pharmacokinetic model[124] Simulate the physiological response of a T1D patient considering recirculation dynamics Closed-loop insulin delivery algorithms Equations and parameters available Evaluate the overhead of runtime as an Android service [125] Medtronic Virtual Patient (MVP) model[126] Explicitly describe intraday variation in metabolic parameters Closed-loop insulin delivery algorithms Equations available Compare the performance of nonlinear filters in online drift detection of CGM[90][50][39] Ackerman’s Linear model [127] Simulate glucose regulation system which is an abstraction of reality Simulate blood glucose control Equations available [128] Simulate and detect replay attack[6] Bergman Minimal Model [128] Describes the dynamics of the glucose uptake after an external stimulus in humans with a simplistic structure Simulate blood glucose control Equations available Evaluate homomorphic encryption [77] Other Mathematical models for T1D [129] Providing a comprehensive overview of the modelling of the biological processes of APS Controller algorithm OpenAPS [130] Implements a controller component of APS in JavaScript Controller’s function Open source Evaluate specification mining technique [86] Other control strategies [129] Providing a comprehensive overview of available control strategies of APS Closed-loop testbed Closed-loop APS testbed [131] Made up of two APS simulators, UVA/Padova and Glucosym, and a fault injection engine APS simulation and faults data generation Open-source Evaluate runtime monitoring unsafe control commands [101][132] MATLAB Simulink model [133] Simulate a CGMS infusion pump for blood glucose regulation APS simulation Open-source Design attacks to compromise sensors or controllers [40] Hardware-in-the-Loop Simulation [115] Simulation of AP using the Bergman Minimal Model (BMM) APS simulation Open-source Simulated the responses of template attack and DoS attack Hardware-in-the-Loop Simulation [116] An low-cost HIL platform for testing control strategies for AP APS simulation Open-source Simulated the PID and linear quadratic regulator (LQR) controllers Datasets OpenHumans [134] Collected by #OpenAPS project, donated by its users BG reading, device function, patient activity Apply for access JCHR - DCLP3 [135] An international diabetes closed-loop trial dataset recorded 6-months from clinical trials with 168 patients BG readings, insulin doses, and other relevant data Publicly-available Test the simulation performance of APS testbed [131] D1NAMO [136] 20 healthy subjects and 9 patients with type-1 diabetes in real-life conditions. ECG, breathing, glucose level and annotated food pictures Publicly-available Evaluating glucose prediction model [137][99] DICARDIA diabetic dataset[138] A diabetic patient database to study cardiovascular autonomic neuropathy (CAN) stress ECG is presented Publicly-available Study the correlation of ECG intervals with blood glucose levels [88] REPLACE-BG dataset[139] CGM data and self-monitoring blood glucose of 226 adult participants with T1D for 26 weeks BG data Publicly-available Re-identify CGM data[108] OhioT1DM Dataset[140] 8 weeks’ worth of continuous glucose monitoring, insulin, physiological sensor, and self-reported life event data for each of 12 people with type 1 diabetes BG level prediction Publicly-available Demonstrate the impact of adversarial inputs on the predictions of the targeted ML model [54] WUSTL-EHMS-2020 dataset [141] Collects both the network flow metrics and patients’ biometrics, labelled as normal and attack instance Network flow, biometrics Publicly-available [142] Detect man-in-the-middle attacks [105]
Table 10: Available resources and utilization in research.
6.1 Hardware Simulation
Hardware simulations offer detailed insights into physical aspects such as communication, energy consumption and heat dissipation, which are crucial for assessing the real-world application of security measures. These simulations focus on evaluating system-level vulnerabilities and performance for defence at the communication and external levels, without involving human participants.
Commercialized diabetes medical devices are key resources for hardware simulation. These devices serve a dual purpose: demonstrating cybersecurity weaknesses and validating defence mechanisms. For example, Li et al. [12] and Marin et al. [48] showed that radio communications in commercial AID systems can be intercepted and compromised using a USRP. Conversely, Zhang et al. [5] used a USRP as the safety monitor for wireless AID systems transactions, and Lazaro et al. [70] tested a communication protocol by combining a commercial CGM, insulin pump, and control algorithm.
Some studies focus on improving the hardware or software of AID systems to enhance security and validate the efficiency of their methods. These studies often use more general and flexible hardware boards to replicate the system. For instance, Strydis et al. [60] used Biostator II, an ultra-low-power ASIC controller, to evaluate the energy efficiency in the proposed security communication protocol. Zhao et al. [79] simulated on an Arduino Uno-R3 microcontroller to assess a visible light access control channel. The Arduino platform was also employed for hardware-in-the-loop simulation to test controller algorithms [116] and simulate the responses of cyber-attacks [115] in AID systems. Prematilake et al. [87] implemented a safety-enhanced insulin pump system on the EFM32WG to monitor both external and internal states. Kim et al. [37] tested an AID systems-specific secure communication protocol using an Arduino Nano 33 for the insulin pump, an nRF52840 kit for the CGM, and a Galaxy S20 for the controller. Ahmed et al.[117] developed a hardware emulation platform using an ESP8266 kit to emulate faults, attacks, and test countermeasures for evaluating closed-loop control systems.
In summary, by leveraging commercial medical devices and custom hardware setups, researchers can identify cybersecurity weaknesses and assess the real-world impact of security measures on AID systems performance.
6.2 Software Simulation
Software simulation, also known as in silico trials, plays a crucial role in evaluating closed-loop insulin delivery systems by mimicking patient responses to various controller strategies and potential cyberattacks. This approach enables researchers to simulate how attacks targeting CGM or insulin delivery commands affect the human body and assess the effectiveness of defence mechanisms at the intrusion detection and control strategies evaluation level.
Accurately simulating the human body’s response in closed-loop insulin delivery systems is crucial for effective pre-clinical testing. One widely used model is the Padova/UVA T1D simulator, an FDA-approved nonlinear simulator for pre-clinical testing [95][96][34][110][59]. It captures interpatient variability by simulating 100 virtual subjects covering children, adolescents, and adults with varying physiological profiles. Since its 2008 release[118], it has evolved to model patient physiology changes over time in 2013 [120] and included intraday insulin sensitivity variations, time-varying therapy parameters, and a ”dawn phenomenon” model in 2018 [121]. Another simulator named Glucosym supports online simulations with 10 real patient profiles, aiding attack simulations and defence evaluations [55].
In addition to these packed simulators, pharmacokinetic models provide a more flexible and transparent approach to simulating physiological responses. These models use mathematical equations and personalized parameters to model T1D patient physiology with greater clarity. For example, Bagade et al. [125] utilize a hybrid physiologically-based pharmacokinetic model [124] to assess the computational overhead of controller algorithms running as Android services. These models are also used in physiologic model-based intrusion detection systems to make predictions [90][50][39].
Controller algorithms are developed to decide insulin dose according to related data to keep the patient’s glucose level within a target range. Among these, OpenAID systems is a notable open-source algorithm that enables T1D patients to use basic AID systems technology for self-management [130]. Other widely used control strategies include Proportional-Integral-Derivative (PID) controllers, Model Predictive Controllers (MPC), fuzzy logic controllers and other nonlinear control strategies. These algorithms provide diverse approaches to insulin management, offering various advantages in terms of responsiveness and adaptability, as detailed in [129], which provides a comprehensive overview of available AID systems control strategies.
To further enhance simulation, some studies integrate T1D patient simulators with controller algorithms to create closed-loop AID systems testbeds. For example, Zhou et al. [131] developed a closed-loop testbed incorporating the UVA/Padova simulator with Basal-Bolus and PID controllers and the Glucosym simulator with a PID-based OpenAID systems controller. They also provide a fault injection engine and a labelled dataset of simulated safe and hazardous records. Another available source is a MATLAB Simulink model testbed, which simulates and integrates a continuous glucose monitor, an insulin pump, and a PID controller [133]. These integrated testbeds facilitate a more convenient and thorough evaluation of AID systems security monitoring systems.
6.3 Real-world datasets
Real-world datasets are important resources for evaluating the effectiveness of cyberattack defence systems in automated insulin delivery systems. Although they often lack explicit attack data, they are valuable for assessing some specific functions like patient modelling and prediction and performance metrics of proposed security mechanisms.
Several datasets provide rich data for continuous glucose monitoring, both from free-living settings and clinical trials. Datasets like JCHR - DCLP3[135], REPLACE-BG [139], D1NAMO[136], and OhioT1DM Dataset [140] provide continuous glucose monitoring data that could be utilized for experiments involving glucose level prediction [99][137] and re-identification [108]. JCHR - DCLP3 dataset [135] was employed by Zhou et al. [131] to test the performance of their proposed AID systems simulation testbed, while Elnawawy et al. [54] used the OhioT1DM Dataset [140] to assess the impact of adversarial inputs on machine learning model predictions.
In addition to glucose monitoring data, some datasets provide supplementary biometric and context information. For example, the DICARDIA dataset [138] contains ECG data, which Panda et al. [88] utilized to develop redundant safety monitors by analyzing the correlation between ECG intervals and blood glucose levels. The WUSTL-EHMS-2020 dataset [141], though not diabetes-specific, offers network flow metrics and biometric data, proving useful for evaluating defence systems that monitor both biometrics and network anomalies, such as detecting man-in-the-middle attacks by Judith et al. [105]. The OpenHumans dataset [134], collected and shared by the OpenAID systems community, includes patient profiles, glucose levels, infusion doses, and ample activity records. Additionally, AID systems-faulty data, generated by the testbed described by Zhou et al.[131] comprises synthetic data from 17,640 closed-loop simulations, including both safe and hazardous data traces from 20 patient profiles, providing a comprehensive source for evaluating fault detection and reaction mechanisms.
7 Key Findings
7.1 RQ1: What potential attack vectors and risks do AID systems face?
The identified attack vectors in Table 5 primarily aim to compromise confidentiality to intercept patient private data or integrity and availability to impact normal insulin delivery, thereby endangering patient health.
Confidential attacks target the wireless communication between the continuous glucose monitor (CGM), controller, and insulin pump, as well as data storage platforms. Attacks on wireless communication will exploit the absence of standardized encryption, enabling eavesdropping using off-the-shelf radio interception toolkits. Notably, even patients themselves might attempt to intercept their data to gain enhanced control over their treatment.
Attacks on availability affect the system’s normal function. For instance, DoS attacks could disrupt transmissions, such as CGM sensor data or controller commands, or potentially drain the system’s battery by repeatedly sending requests. They may be executed by cyber-physical attacks, requiring a moderate level of knowledge regarding the relevant tools and communication protocols, without necessarily understanding the underlying control processes of the system.
As a closed-loop control system, the insulin dosage can also be manipulated by compromising the integrity of any component within the loop. This includes altering CGM sensor data, food state information, controller data, controller algorithms, controller commands, and the insulin pump driver. Replay attacks are easier to launch without requiring the knowledge of the communication protocol. Random data injection attacks and bias data injection attacks demand higher adversarial capabilities, as they involve reverse-engineering the communication protocol and fortifying the manipulated data to be accepted by the system. Nevertheless, such attacks have been successfully carried out on commercial AID systems products [12][48].
A more refined attack method involves leveraging the knowledge of the system’s control process to manipulate the system into a harmful but acceptable state. Subtlety in this context can be interpreted from two perspectives. The first is personalised attacks, where a specific dosage may be harmless to the general population but can cause significant harm to individual patients under certain conditions. The second perspective relates to ML/DL models used in controller algorithms. These models, which are often black-box in nature, allow attacks that introduce slight changes in input data to make a significant impact on the model[53]. Such attacks require advanced knowledge of the patient and the system. However, it is possible to be obtained by intercepting private patient information. Detailed computational algorithms and model weights might be found in the product’s technical documentation or be estimated by intercepting input-output pairs.
7.2 RQ2: How do existing defence mechanisms address these risks?
Existing proposed mechanisms protect AID systems at different layers, as shown in Table 7.
The first layer focuses on securing communication to prevent data interception or manipulation. Techniques such as authentication, cryptography, and secure communication protocols can be employed to safeguard wireless communication channels from external interference. Another approach is to use proximity communication channels to avoid in-the-air communication, such as body-coupled communication or visible-light communication. In the future, there may also be innovations involving materials that control insulin injection via tubing [143].
The second layer involves the use of intrusion detection systems. At this stage, data or system functions may already have been compromised. IDS are deployed to monitor system activity and detect any predefined threats or deviations from normal behaviour. These methods can address known threats, unknown attacks that deviate from normal operations, and unusual data patterns.
The third layer involves assessing control strategies. In cases where an attacker has successfully altered data, control algorithms, or system configurations, this layer ensures that any insulin injection command issued accords with predefined thresholds or the patient’s normal injection patterns. This layer acts as a final check to validate the accuracy and safety of the system’s operation before injection.
Redundant hardware or software functions as additional protection without being incorporated into the AID systems. This approach offers two main advantages. First, the protection model does not consume the system’s battery, memory, or computational resources. Second, since the protection model operates independently of the insulin infusion system, it does not require recertification. This allows for the faster application of security patches in response to new threats. These methods include using extra hardware or software to monitor communication traffic or content, as well as incorporating additional sensors to detect deviations.
7.3 RQ3: How practical and robust are these solutions in protecting AID systems from cybersecurity threats?
As outlined in Table 6, most threats have corresponding defence strategies that can be applied to protect AID systems. However, the effectiveness of these strategies in providing robust protection without exceeding resource requirements depends. Additionally, effective methods to defend against advanced process-aware attacks are lacking. Achieving robust and practical protection for AID systems requires consideration of several key factors:
  1. 1.
    Overall security: While current defence methods target specific threats and operate within different protection layers, a reliable and robust AID system solution should address all potential threats and respond to emergencies in a timely manner. Fully securing an AID system requires a comprehensive approach that integrates these layers.
  2. 2.
    Computational overhead: Despite many defence methods considering the computational overhead in design, the cumulative and combined overhead of the application has not been thoroughly evaluated. Certain techniques, such as homomorphic encryption and machine learning-based IDS, have significant computational demands on their own, which could impact the system’s overall performance when integrated into AID systems.
  3. 3.
    Dynamic analysis: Given that AID systems devices interact directly with the human body, the complexity of device functions, individual patient differences, and fluctuating health conditions make it difficult to define fixed behaviours for detecting malicious activity. Most defence methods rely on fixed definitions or models that may not adapt to personalized features, evolving patient conditions, or changing attack vectors, which could impair their ability to maintain optimal performance over time.
  4. 4.
    Free-living use: Many defence strategies are based on unrealistic assumptions, such as patients adhering to a rigid routine, having a fixed injection pattern, being attacked only during specific times (e.g., at night), or not engaging in meals or exercise. These assumptions do not hold in the typical free-living context of AID systems use. Meal taking, physical activity, and other personal factors all affect glucose levels and insulin requirements. Failing to account for these variables may result in frequent false alarms or ineffective detection of threats.
8 Open Research Challenges and Future Directions
Recent advancements in the security of AID systems have bolstered their resilience, particularly in wireless communication, sensor data integrity, and control mechanisms. However, the pursuit of highly reliable and secure AID systems continues to still face several significant challenges, many of which remain unresolved.
8.1 Resource Constraints in Wearable AID systems Devices
AID systems devices are inherently limited in terms of processing power, memory, and battery life, making the integration of resource-intensive security measures particularly challenging. Many defence mechanisms research only tested risk identification performance in silico data without considering resource consumption. Implementing such advanced security features will introduce computational burdens that can compromise the performance and functionality of AID systems. Balancing security performance with the system’s limited resources is a pressing concern, especially as AID systems must remain lightweight enough to be wearable and function effectively without draining power or degrading response time.
8.2 Variability Across Patients
A major challenge in securing AID systems lies in the significant variability between patients, both within and across different individuals (intra- and inter-patient variability). Each patient’s physiological responses, lifestyle, and treatment regimen are unique, making it difficult to develop universal security models that adapt to these personalized characteristics. Personalized cyberattacks targeting these patient-specific variabilities pose a considerable threat, and current security systems are often rigid, failing to accommodate such nuances. Future AID systems security models need to dynamically account for patient heterogeneity, while still offering robust protection against diverse cyber threats.
8.3 Escalating Complexity of Cyberattacks
The evolving nature of cyberattacks presents another formidable challenge for AID systems security. Attackers frequently develop sophisticated, multi-vector strategies that can exploit multiple vulnerabilities in the system, from wireless communication protocols to insulin control algorithms. While existing defences often focus on countering specific attack types and assume other parts are intact, these solutions will fall short when faced with multi-faceted or cross-layer attacks that target different components of AID systems simultaneously. Future defence systems must be capable of addressing these complex and hybrid attack vectors.
8.4 Comprehensive Modelling of Insulin Infusion Patterns
Accurately predicting insulin needs remains a challenge, particularly in response to dynamic physiological states such as meal intake, physical activity, and stress. Current insulin infusion models struggle to account for the full range of variables affecting glucose levels. There is a need for comprehensive models that integrate real-time data to predict and adapt to patient-specific insulin requirements. Improving the accuracy of these models not only facilitates the development of fully closed-loop insulin control systems but also plays a critical role in defending against cyberattacks that manipulate injected doses within insulin delivery systems.
8.5 Standardization of AID systems Cybersecurity Frameworks
One of the most critical gAID systems in AID systems security is the lack of standardized frameworks for assessing the conformity and robustness of cybersecurity implementations. Manufacturers, regulators, and healthcare providers need to collaborate on establishing uniform standards to ensure that all AID systems devices undergo rigorous security testing. These standards should provide confidence in the device’s ability to autonomously detect and respond to threats in real time. Developing a standardized assessment framework would enable consistent evaluation of AID systems’ security, propelling the development of highly reliable medical devices.
8.6 Trustworthy Defense Mechanisms
While securing AID systems is crucial, equally important is ensuring that these defence mechanisms are explainable and preserve patient privacy. Recent data-driven cybersecurity solutions often operate as black boxes offering little to no insight into how decisions are made, making it difficult for users or healthcare professionals to interpret or intervene. Furthermore, the collaborative nature of future AID systems networks—where data is shared between devices to improve security—presents additional privacy concerns. Effective knowledge sharing without breaching patient confidentiality remains a delicate balance, especially when dealing with sensitive patient data.
9 Conclusion
This review has provided a comprehensive analysis of the current state of AID systems security, exploring existing attack vectors, defence mechanisms, and available research resources such as simulators and real-world datasets. Several critical challenges were highlighted for the need for further investigation. In conclusion, while substantial progress has been made in securing AID systems, continued innovation is required to develop adaptive, resource-efficient, and patient-centric defence mechanisms. The path forward demands collaborative efforts from researchers, manufacturers, regulators, and healthcare providers to establish standardized frameworks and ensure that AID systems are effective and secure in protecting patients from evolving cyber threats. By addressing these challenges, AID systems can be trusted and widely adopted by diabetes patients to improve their quality of life, while maintaining the highest standards of safety and security.
Acknowledgment
This work was supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2, under Grant MOE-T2EP20121-0008.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to improve the readability and language of the manuscript. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.